# Evolution and rotation of optically bound matter in focused laser beams

### v1

@@ Abstract
Optical matter (OM), composed of optically bound metallic particles, can be rotated by transferring the spin angular momentum (SAM) of chiral light to the assembly. Rotating OM is a promising platform for optical micromachines, with potential applications in plasmofluidics and soft robotics. Understanding the dynamic states of such Brownian, micro-mechanical systems is a relevant issue. One key problem is understanding kinetic jamming and clogging. Studies of driven multiparticle systems have revealed that under suboptimal driving, the systems can stop moving, showing jamming transitions. It is important to identify dynamic regimes where crowding competes with driving and is susceptible to jamming in the context of optical micromachines. Through experiments supported by numerical simulations, we reveal assemblies with well-defined hexagonal or triangular symmetry that efficiently harness the SAM of incident chiral light, resulting in stable rotation. However, as the plasmonic-particle assembly grows and its dimensions approach the beam waist, new particles can disrupt this order. This causes a transition to a ‘fluid-like’ state with less-defined symmetry, correlated with a significant reduction in transferred torque, causing rotation to stagnate or cease. We suggest this behaviour is analogous to a rotational jamming transition, where the rotational motion is arrested. Our findings establish a clear relationship between the structural symmetry of the OM assembly and its ability to harness SAM, providing new insights into controlling chiral light-matter interactions and offering a novel platform for studying jamming transitions.
@@
[…]
Present Address: Faculty of Engineering and Natural Sciences, Tampere University, FI-33014, Tampere, Finland Unknown University] Department of Physics, Indian Institute of Science Education and Research Pune, Pune, Maharashtra, 411008, India
@@ 1 Introduction
Optical matter (OM) is a dynamic, non-equilibrium phase of matter composed of particles that self-assemble and organize under continuous illumination from an off-resonant laser.thirunamachandranIntermolecularInteractionsPresence1980, burnsOpticalBinding1989c, mcgloinOpticallyBoundMicroscopic2004a, mohantyOpticalBindingDielectric2004c, grzegorczykStableOpticalTrapping2006c, karasekAnalysisOpticalBinding2006b, ahlawatOpticalTrappingCharged2007, karasekLongRangeOneDimensionalLongitudinal2008c, demergisUltrastrongOpticalBinding2012d, aritaOpticalBindingTwo2018 The particles are held together by optical binding force, which is a laser scattering-induced interaction.dholakiaColloquiumGrippedLight2010c, forbesOpticalBindingNanoparticles2020d, rodriguezOpticalBindingNanoparticle2008b, romeroElectrodynamicMechanismArray2008b, demergisUltrastrongOpticalBinding2012d, kudoOpticalTrappingFormedColloidal2016, kudoSingleLargeAssembly2018, wangOpticallyEvolvedAssembly2016 This interaction creates stable, dynamic configurations in which particles are held together without physical contact, often with interparticle distances that are integer multiples of the laser wavelength.yanGuidingSpatialArrangements2013b, simpsonOpticalBindingNanowires2017b, hanCrossoverPositiveNegative2018a, hanPhaseTransitionSelfStabilization2020c, louisUnconventionalOpticalMatter2024b Optical matter, therefore, represents a unique system for studying self-organization and collective behavior in driven, non-equilibrium systems.hanPhaseTransitionSelfStabilization2020c, hanCrossoverPositiveNegative2018a, qiStableNegativeOptical2022a, taoRotationalDynamicsIndirect2023a, taoGeneralizedOpticalBinding2024c, chenNegativeOpticalTorque2014 Beyond linear momentum, light also possesses spin angular momentum (SAM), which is associated with its polarization state.sharmaSpinHallEffectScattering2018, sharmaOpticalOrbitalAngular2019, paulFocusedLinearlypolarizedlightScattering2021, paulSimultaneousDetectionSpin2022, bliokhSpinOrbitInteractions2015, liawSpinOrbitalRotation2018b, kotlyarSpinorbitOrbitspinConversion2020b, kumarWhenPlasmonicColloids2023 This SAM can be transferred to colloids as well as optical matter assembly, causing them to rotate.stilgoeControlledTransferTransverse2022, changOpticalTorqueExerted1998b, padgettLightTwistIts2000b, bruceInitiatingRevolutionsOptical2021a, frieseOpticallyDrivenMicromachine2001b, toftulOpticalSupertorqueInduced2025, patersonControlledRotationOptically2001b, adachiOrbitalMotionSpherical2007b, zhaoDirectObservationTransfer2009b, lehmuskeroUltrafastSpinningGold2013c, yuRotationDynamicsParticles2015b, shaoLightDrivenRotationPlasmonic2018b, reichertHydrodynamicCouplingTwo2004 The magnitude and direction of the imparted torque are determined by the specific structure of the particle assembly and the chirality of the illuminating laser beam.rodriguezOpticalBindingInfluence2008, brzobohatyExperimentalTheoreticalDetermination2010, sukhovActioReactioOptical2015, haefnerConservativeNonconservativeTorques2009, yanPotentialEnergySurfaces2014a, huangPrimevalOpticalEvolving2022 In assemblies with dynamical states having well-defined rotational symmetry, the scattering of light is predictable. This facilitates an efficient and continuous transfer of angular momentum. The potential application of OM as micromachines, including micro-gears and micro-engines,figliozziDrivenOpticalMatter2017a, parkerOpticalMatterMachines2020a, wangMicroscopicGearedMetamachines2025 and in soft robotics relies on the ability to precisely control these dynamical states.
@@
Understanding the degrees of freedom of such light-driven metallic micromachines has emerged as a relevant problem in optically bound matter. To reveal the physical conditions under which OM can be harnessed as rotational micromachines, it is necessary to interrogate them as micro-mechanical systems subject to Brownian noise. This framework of driven systems with Brownian noise motivates a statistical mechanics viewpoint, where several concepts related to dynamic states have been discussed and studied at micron and sub-micron scales.nagelExperimentalSoftmatterScience2017a Among the many, jamming of soft matter systemsliuJammingNotJust1998 has turned out to be of interest not only as a curious observation, but also with technological implications. Jamming transitions are a universal kinetic arrest phenomenon that has been widely studied in macroscopic granular media, foams, and colloids.chenOpticalManipulationSoft2024, sellerioGlassTransitionAssociated2011, kumarStructuralDynamicProperties2004, zhangJammingTransitionEmulsions2005, ravazzanoUnjammingActiveRotators2020, liuLocalRotationalJamming2023, zhuVariableStiffnessSoft2024, behringerPhysicsJammingGranular2018, ganapathiStructureDeterminesWhere2021, biJammingShear2011, henkesJammingCriticalPhenomenon2005, himanagamanasaDirectMeasurementsGrowing2015, zottlEmergentBehaviorActive2016, tanOddDynamicsLiving2022 In these systems, increasing particle density leads to a disordered, jammed state where motion is arrested. The efficient driving of the systems requires the identification of a suitable parameter space of particle density and driving.cereceda-lopezHydrodynamicInteractionsCan2021
@@
In this report, we experimentally and computationally demonstrate the jamming of rotating OM formed by 400 nm gold nanoparticles, assembled and driven by a SAM-carrying focused beam. Focused laser beams provide a more efficient method for optical manipulation,jonesOpticalTweezersPrinciples2015, volpeRoadmapOpticalTweezers2023b, chandOptothermalRevolutionColloids2025, shuklaSynchronizedMotionGold2025a and owing to their finite extent, a crowding effect is possible where multiple particles are present in a confined region. It is already known that the behavior of optical matter changes fundamentally under focused beams due to the strong intensity gradients.yanPotentialEnergySurfaces2014a, huangPrimevalOpticalEvolving2022a Thus, focused beams provide an excellent platform to study jamming in rotating optical matter. The jamming we observe occurs due to a structural transition, as shown schematically in Figure 1a. As a stable, hexagonally packed assembly of gold nanoparticles grows, the addition of a new particle can disrupt its order, causing a dramatic transition from a rigid, crystalline-like structure to a disordered, ‘fluid-like’ state. This loss of symmetry is directly correlated with a significant reduction in transferred torque, causing rotation of the assembly to stagnate or cease. The observations reported here are similar to jamming transitions. Our system, driven by a continuous optical force rather than thermal motion, offers a unique mesoscopic platform for studying these phenomena.
@@
2 Results and discussion
Building on the discussed framework, we now present the experimental and computational evidence for the rotational arrest phenomenon. Optically bound nanoparticles have been studied under both focused and broad illumination. Broad illumination requires larger input powers to have sufficient intensity in the sample plane, as it depends inversely on the square of the beam waist. Foundational insights have been gained from OM rotation experiments and simulations, considering both broad-beam and plane-wave illuminations that carry spin angular momentum. But the behaviour of OM fundamentally changes under focused beams due to strong intensity gradients, which is the focus of our study. Thus, we demonstrate the evolution of rotating optically bound matter (OM) under conditions of a circularly polarized focused beam, which enables the creation of assemblies using a minimal power of approximately 20 mW. We illustrate the changes due to focusing by experimenting with two different numerical aperture objectives, which make beam spots of 1.5 $\mu m$ and 2 $\mu m$, respectively. We have discussed the theoretical differences in optical binding in plane wave vs focused beam illumination in Supplementary Information S1 and S2.
@@ 2.1 Arrested Rotation of Four Particle Assembly
The observed stopping of rotation in our optical matter assembly can be described as a dynamic arrest phenomenon observed in soft matter.chenOpticalManipulationSoft2024, nagelExperimentalSoftmatterScience2017a, eberleDynamicalArrestTransition2011, reichhardtActiveMatterTransport2014 In this system, the optical torque, which provides the drive or activity, competes with increasing particle number, which introduces a crowding effect. This transition from a rotating, ordered state to a jammed, non-rotating state shares striking parallels with dense active systems, where a transition occurs as the drive competes with the increasing density of active particles. In our case, the addition of a single particle acts as a defect that disrupts the collective symmetry, effectively jamming the system and arresting the motion. This suggests that the interplay between symmetry and torque in light-driven systems can be viewed through the lens of collective dynamics, offering a new perspective on controlling light-matter interactions at the nanoscale.
@@ 2.2 Arrested Rotation of Eight Particle Assembly: Effect of Beam Size
We show representative snapshots of 7P and 8P assembly in Figure 3a. The overall trajectory of all particles is shown in the insets. To quantify the loss of rotation, the mean angular position of the peripheral particles is shown in Figure 3b. The angles increase until the 8th particle joins the assembly, and then stagnate and fluctuate. Finally, to quantify the structural order, we calculate a local hexagonal order parameter.sharmaLargescaleOptothermalAssembly2020, hanPhaseTransitionSelfStabilization2020c This order parameter is defined as:
@@
where i sums over all the N particles which are neighbours of the central particle in the assembly. Since our system only consists of 7 to 8 particles, the $\psi_{6}$ gives information about the sixfold symmetry of the assembly. When the assembly is sixfold symmetric, HOP has a value equal to 1, and decreasing sixfold symmetry implies a decrease in HOP up to 0. This local hexagonal order parameter is plotted for all the frames of the Supplementary Video 2 in Figure 3c. The order parameters for 7P and 8P times are colored differently for contrast. The order parameter reduces for the 8P assembly. We plot the order parameter distribution histogram in the same figure. The histogram shows that for 7P assembly, the HOP is more skewed to 1, which implies a hexagonal packing, whereas the 8P assembly has a peak near 0.4. Typically, particles in a fluid have a HOP value around 0.4.hanPhaseTransitionSelfStabilization2020c Thus, we refer to this state as a ‘fluid-like’ state.
@@
As noted previously, this loss of rotation is akin to more widely observed arresting phenomena. But in the case of eight particles, we can see the role of a hexagonal order more clearly compared to the four-particle case discussed in the previous section. This transition from an ordered, crystalline-like state ($\psi_{6}\approx 1$) to a disordered, amorphous state ($\psi_{6}\approx 0.4$) can be described as a jamming transition.libalColloidalLatticeShearing2013, senbilObservationStronglyHeterogeneous2019, kordaEvolutionColloidalCritical2002, caoNumericalStudyDynamics2003 Our observation represents a rotational jamming transition, a form of universal kinetic arrest phenomenon where the rotational motion of the system ceases.robertson-andersonOpticalTweezersMicrorheology2018, cereceda-lopezHydrodynamicInteractionsCan2021, lipsHydrodynamicInteractionsHinder2022, cereceda-lopezExcludedVolumeInduces2024, saitoChangeCollectiveMotion2018a, yanRotatingCrystalsMagnetic2015, stoopCloggingJammingColloidal2018 In this dynamic, light-driven system, the seven-particle assembly maintains its crystalline-like hexagonal structure, enabling stable rotation. However, the addition of the eighth particle disrupts this order, causing the system to exceed a critical packing threshold. The eight-particle assembly, with its disordered structure, represents a kinetically arrested transient phase. This state is unable to efficiently convert light’s spin angular momentum into collective orbital motion, causing the rotational drive to be overcome by the loss of symmetry.
@@
To further validate our findings and explore the effect of different parameters, we investigated assemblies of 250 nm gold nanoparticles (AuNPs) using a 785 nm laser, shown in Supplementary Video 3 and discussed in Supplementary Information S3. While previous literatureqiStableNegativeOptical2022a has shown that particles ranging from 100 nm to 350 nm can form stable, rotating assemblies, our study demonstrates that 400 nm AuNPs ($\approx 0.5\lambda_{m}$) can also form rotating optical matter. However, the symmetry transition and subsequent rotation arrest are highly dependent on the particle size, laser wavelength, and beam profile. Our experiments confirm that varying these parameters leads to different magnitudes of scattering forces, which in turn give rise to non-hexagonal states, as shown in Figure 4.

### v2

# Rotational Jamming of Plasmonic Optical Matter Driven by Chiral Light

@@ Abstract
Plasmonic Optical matter (OM), composed of optically bound metallic particles, can be rotated by transferring the spin angular momentum (SAM) of chiral light to the assembly. Rotating OM is a promising platform for optical micromachines, with potential applications in plasmofluidics and soft robotics. Understanding the dynamic states of such Brownian, micro-mechanical systems is a relevant issue. One key problem is understanding kinetic jamming and clogging. Studies of driven multiparticle systems have revealed that under suboptimal driving, the systems can stop moving, showing jamming transitions. It is important to identify dynamic regimes where crowding competes with driving and is susceptible to jamming in the context of optical micromachines. Through experiments supported by numerical simulations, we reveal assemblies with well-defined hexagonal or triangular symmetry that efficiently harness the SAM of incident chiral light, resulting in stable rotation. However, as the plasmonic-particle assembly grows and its dimensions approach the beam waist, new particles can disrupt this order. This causes a transition to a ‘fluid-like’ state with less-defined symmetry, correlated with a significant reduction in transferred torque, causing rotation to stagnate or cease. We suggest this behaviour is analogous to a rotational jamming transition, where the rotational motion is arrested. Our findings establish a clear relationship between the structural symmetry of the OM assembly and its ability to harness SAM, providing new insights into controlling chiral light-matter interactions and offering a novel platform for studying jamming transitions.
@@ 1 Introduction
Optical matter (OM) is a dynamic, non-equilibrium phase of matter composed of particles that self-assemble and organize under continuous illumination from an off-resonant laser.thirunamachandran_intermolecular_1980, burns_optical_1989, mcgloin_optically_2004, mohanty_optical_2004, grzegorczyk_stable_2006, karasek_analysis_2006, ahlawat_optical_2007, karasek_long-range_2008, demergis_ultrastrong_2012, arita_optical_2018 The particles are held together by optical binding force, which is a laser scattering-induced interaction.dholakia_colloquium_2010, forbes_optical_2020, rodriguez_optical_2008, romero_electrodynamic_2008, demergis_ultrastrong_2012, kudo_optical_2016, kudo_single_2018, wang_optically_2016 This interaction creates stable, dynamic configurations in which particles are held together without physical contact, often with interparticle distances that are integer multiples of the laser wavelength.yan_guiding_2013, simpson_optical_2017, chen_negative_2014, qi_stable_2022, yan_potential_2014, huang_primeval_2022, mirzaei-ghormishNonlinearOpticalBinding2025b Optical matter, therefore, represents a unique system for studying self-organizationhan_crossover_2018, han_phase_2020, louis_unconventional_2024, tao_generalized_2024, tao_rotational_2023 and collective behavior in driven, non-equilibrium systems. Beyond linear momentum, light also possesses spin angular momentum (SAM), which is associated with its polarization state.sharma_spin-hall_2018, sharma_optical_2019, paul_focused_2021, paul_simultaneous_2022, bliokh_spinorbit_2015, liaw_spin_2018, kotlyar_spin-orbit_2020, kumar_when_2023 This SAM can be transferred to colloids as well as optical matter assembly, causing them to rotate.stilgoe_controlled_2022, chang_optical_1998, padgett_light_2000, bruce_initiating_2021, friese_optically_2001, toftul_optical_2025, paterson_controlled_2001, adachi_orbital_2007, zhao_direct_2009, lehmuskero_ultrafast_2013, yu_rotation_2015, shao_light-driven_2018, reichert_hydrodynamic_2004 The magnitude and direction of the imparted torque are determined by the specific structure of the particle assembly and the chirality of the illuminating laser beam.rodriguez_optical_2008-1, brzobohaty_experimental_2010, sukhov_actio_2015, haefner_conservative_2009, han_crossover_2018, genet_chiral_2022 In assemblies with dynamical states having well-defined rotational symmetry, the scattering of light is predictable. This facilitates an efficient and continuous transfer of angular momentum. The potential application of OM as micromachines, including micro-gears and micro-engines,figliozzi_driven_2017, parker_optical_2020, wang_microscopic_2025 and in soft robotics relies on the ability to precisely control these dynamical states.
@@
Understanding the degrees of freedom of such light-driven metallic micromachines has emerged as a relevant problem in optically bound matter. To reveal the physical conditions under which OM can be harnessed as rotational micromachines, it is necessary to interrogate them as micro-mechanical systems subject to Brownian noise. This framework of driven systems with Brownian noise motivates a statistical mechanics viewpoint, where several concepts related to dynamic states have been discussed and studied at micron and sub-micron scales.nagel_experimental_2017, martinezBrownianCarnotEngine2016 Among the many, jamming of soft matter systemsliu_jamming_1998 has turned out to be of interest not only as a curious observation, but also with technological implications. Jamming transitions are a universal kinetic arrest phenomenon that has been widely studied in macroscopic granular media, foams, and colloids.chen_optical_2024, sellerio_glass_2011, kumar_structural_2004, zhang_jamming_2005, ravazzano_unjamming_2020, liu_local_2023, zhu_variable_2024, behringer_physics_2018, ganapathi_structure_2021, bi_jamming_2011, henkes_jamming_2005, hima_nagamanasa_direct_2015, zottl_emergent_2016, tan_odd_2022, zaccone_complete_2025 In these systems, increasing particle density leads to a disordered, jammed state where motion is arrested. The efficient driving of the systems requires the identification of a suitable parameter space of particle density and driving.cereceda-lopez_hydrodynamic_2021
@@
In this report, we experimentally and computationally demonstrate the jamming of rotating OM formed by 400 nm gold nanoparticles, assembled and driven by a SAM-carrying focused beam. Focused laser beams provide a more efficient method for optical manipulation,jones_optical_2015, volpe_roadmap_2023, chand_optothermal_2025, shukla_synchronized_2025, tsujiThermoosmoticSlipFlows2023 and owing to their finite extent, a crowding effect is possible where multiple particles are present in a confined region. It is already known that the behavior of optical matter changes fundamentally under focused beams due to the strong intensity gradients.yan_potential_2014, huang_primeval_2022 Thus, focused beams provide an excellent platform to study jamming in rotating optical matter. The jamming we observe occurs due to a structural transition, as shown schematically in Figure 1a. As a stable, hexagonally packed assembly of gold nanoparticles grows, the addition of a new particle can disrupt its order, causing a dramatic transition from a rigid, crystalline-like structure to a disordered, ‘fluid-like’ state. This loss of symmetry is directly correlated with a significant reduction in transferred torque, causing rotation of the assembly to stagnate or cease. The observations reported here are similar to jamming transitions. Our system, driven by a continuous optical force rather than thermal motion, offers a unique mesoscopic platform for studying these phenomena.
@@
2 Results
Building on the discussed framework, we now present the experimental and computational evidence for the rotational arrest phenomenon. Optically bound nanoparticles have been studied under both focused and broad illumination. Broad illumination requires larger input powers to have sufficient intensity in the sample plane, as it depends inversely on the square of the beam waist. Foundational insights have been gained from OM rotation experiments and simulations, considering both broad-beam and plane-wave illuminations that carry spin angular momentum. But the behaviour of OM fundamentally changes under focused beams due to strong intensity gradients, which is the focus of our study. Thus, we demonstrate the evolution of rotating optically bound matter (OM) under conditions of a circularly polarized focused beam, which enables the creation of assemblies using a minimal power of approximately 20 mW. We illustrate the changes due to focusing by experimenting with two different numerical aperture objectives, which make beam spots of 1.5 $\mu m$ and 2 $\mu m$, respectively. We have discussed the theoretical differences in optical binding in plane wave vs focused beam illumination in Supplementary Information S1 and S2supple.
@@ 2.1 Arrested Rotation of Four Particle Assembly
The observed stopping of rotation in our optical matter assembly can be described as a dynamic arrest phenomenon observed in soft matter.chen_optical_2024, nagel_experimental_2017, eberle_dynamical_2011, reichhardt_active_2014 In this system, the optical torque, which provides the drive or activity, competes with increasing particle number, which introduces a crowding effect. This transition from a rotating, ordered state to a jammed, non-rotating state shares striking parallels with dense active systems, where a transition occurs as the drive competes with the increasing density of active particles. In our case, the addition of a single particle acts as a defect that disrupts the collective symmetry, effectively jamming the system and arresting the motion. This suggests that the interplay between symmetry and torque in light-driven systems can be viewed through the lens of collective dynamics, offering a new perspective on controlling light-matter interactions at the nanoscale.
@@ 2.2 Arrested Rotation of Eight Particle Assembly: Effect of Beam Size
We show representative snapshots of 7P and 8P assembly in Figure 3a. The overall trajectory of all particles is shown in the insets. To quantify the loss of rotation, the mean angular position of the peripheral particles is shown in Figure 3b. The angles increase until the 8th particle joins the assembly, and then stagnate and fluctuate. Finally, to quantify the structural order, we calculate a local hexagonal order parameter.sharma_large-scale_2020, han_phase_2020 This order parameter is defined as:
@@
where i sums over all the N particles which are neighbours of the central particle in the assembly. Since our system only consists of 7 to 8 particles, the $\psi_{6}$ gives information about the sixfold symmetry of the assembly. When the assembly is sixfold symmetric, HOP has a value equal to 1, and decreasing sixfold symmetry implies a decrease in HOP up to 0. This local hexagonal order parameter is plotted for all the frames of the Supplementary Video 2 in Figure 3c. The order parameters for 7P and 8P times are colored differently for contrast. The order parameter reduces for the 8P assembly. We plot the order parameter distribution histogram in the same figure. The histogram shows that for 7P assembly, the HOP is more skewed to 1, which implies a hexagonal packing, whereas the 8P assembly has a peak near 0.4. Typically, particles in a fluid have a HOP value around 0.4.han_phase_2020 Thus, we refer to this state as a ‘fluid-like’ state.
@@
As noted previously, this loss of rotation is akin to more widely observed arresting phenomena. But in the case of eight particles, we can see the role of a hexagonal order more clearly compared to the four-particle case discussed in the previous section. This transition from an ordered, crystalline-like state ($\psi_{6}\approx 1$) to a disordered, amorphous state ($\psi_{6}\approx 0.4$) can be described as a jamming transition.libal_colloidal_2013, senbil_observation_2019, korda_evolution_2002, cao_numerical_2003 Our observation represents a rotational jamming transition, a form of universal kinetic arrest phenomenon where the rotational motion of the system ceases.robertson-anderson_optical_2018, cereceda-lopez_hydrodynamic_2021, lips_hydrodynamic_2022, cereceda-lopez_excluded_2024, saito_change_2018, yan_rotating_2014, stoop_clogging_2018 In this dynamic, light-driven system, the seven-particle assembly maintains its crystalline-like hexagonal structure, enabling stable rotation. However, the addition of the eighth particle disrupts this order, causing the system to exceed a critical packing threshold. The eight-particle assembly, with its disordered structure, represents a kinetically arrested transient phase. This state is unable to efficiently convert light’s spin angular momentum into collective orbital motion, causing the rotational drive to be overcome by the loss of symmetry.
@@
To further validate our findings and explore the effect of different parameters, we investigated assemblies of 250 nm gold nanoparticles (AuNPs) using a 785 nm laser, shown in Supplementary Video 3 and discussed in Supplementary Information S3.supple While previous literatureqi_stable_2022 has shown that particles ranging from 100 nm to 350 nm can form stable, rotating assemblies, our study demonstrates that 400 nm AuNPs ($\approx 0.5\lambda_{m}$) can also form rotating optical matter. However, the symmetry transition and subsequent rotation arrest are highly dependent on the particle size, laser wavelength, and beam profile. Our experiments confirm that varying these parameters leads to different magnitudes of scattering forces, which in turn give rise to non-hexagonal states, as shown in Figure 4.